OpenStudy (anonymous):

how do i determine whether 2 lines are parallel, perpendicular, or niether?

OpenStudy (anonymous):

perpendicular lines intersect

OpenStudy (anonymous):

parallel lines never meet

OpenStudy (anonymous):

i need to know that from the points (-3,-2) (9,1) (3,6) and (5,-2)

OpenStudy (ashleyisakitty):

Graph them on a graphing calculator, or WolframAlpha, then follow the rules that @dmart180 explained.

OpenStudy (anonymous):

ok

OpenStudy (ashleyisakitty):



OpenStudy (ashleyisakitty):



OpenStudy (anonymous):

don't listen to @Ashleyisakitty, drawing are not been used to determine if two lines are perpendicular or parallel because they lack serious rigorousity

OpenStudy (anonymous):

what u should do for parallelism test is to find if the a value in their y=ax+b is the same

OpenStudy (ashleyisakitty):

Its a general example of what to look for, you can also follow the equation.

OpenStudy (arabpride):

~>If two equations have the SAME SLOPE, which is represented by the "m" variable in the form shown above, then these two equations are PARALLEL [ex: y = 3x + 7 is parallel to y = 3x + 1] ~>If two equations have an OPPOSITE RECIPROCAL SLOPE, then they are PERPENDICULAR to one another [ex: y = -1/2x + 3 is perpendicular to y = 2x + 9] ~>If two equations have neither of these relationships, then your answer would be "neither". Make sense? :3